# Revisit What You See: Disclose Language Prior in Vision Tokens for Efficient Guided Decoding of LVLMs

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across various multimodal tasks by integrating visual perception with language understanding. However, conventional decoding strategies of LVLMs often fail to successfully utilize visual information, leading to visually ungrounded responses. While various approaches have been proposed to address this limitation, they typically require additional training, multi-step inference procedures, or external model dependencies. This paper introduces ReVisiT, a simple yet effective decoding method that references vision tokens to guide the text generation process in LVLMs. Our approach leverages the semantic information embedded within vision tokens by projecting them into the text token distribution space, and dynamically selecting the most relevant vision token at each decoding step through constrained divergence minimization. This selected vision token is then used to refine the output distribution to better incorporate visual semantics. Experiments on three LVLM hallucination benchmarks with two recent LVLMs demonstrate that ReVisiT consistently enhances visual grounding with minimal computational overhead. Moreover, our method achieves competitive or superior results relative to state-of-the-art baselines while reducing computational costs for up to 2×2$\times$2 ×. [^1] Code is available at https://github.com/bscho333/ReVisiT.
@@ 1 Introduction
With the recent success of Large Language Models (LLMs; [1, 2, 3, 4]), Large Vision-Language Models (LVLMs; [5, 6, 7, 8, 9, 10, 11, 12, 13]) have emerged as powerful multimodal architecture that integrate visual perception with language understanding and text generation. LVLMs typically incorporate visual information by encoding images into embeddings using a vision encoder, then mapping them to vision tokens via a cross-modality projector aligned with LLM decoder’s linguistic embedding space. Then, LVLMs utilize these vision tokens alongside text tokens during the text generation in autoregressive manner. These models have demonstrated impressive performance on foundational vision-language tasks such as Visual Question Answering (VQA; [14, 15, 16]) and image captioning [17, 18, 19], while also showing promising results in more complex reasoning tasks, e.g., multimodal conversational agents [20]. Despite these advances, the decoding process of LVLMs largely follows conventional LLM decoding strategies without specific consideration for vision tokens. This insufficient extraction and utilization of visual information often result in visually ungrounded responses [21, 22, 23, 24, 25, 26, 27, 28, 29].
Prior efforts to address this limitation can be broadly categorized based on when and how additional visual information is incorporated relative to the decoding process. Specifically, we distinguish between (1) pre-alignment, (2) post-alignment, and (3) intra-alignment. Pre-alignment methods aim to enhance visual information before the decoding phase, typically by re-designing model architectures [30] or developing specialized visual tokenization techniques [31, 32], but they generally require additional training and model modification. Post-alignment strategies revise the generated outputs through external models, such as object detectors or proprietary models [21, 22], inevitably introducing complexity and external dependencies. Due to these inherent overheads of pre- and post-alignment methods, researchers have increasingly focused on intra-alignment approaches, which modify attention mechanisms or output distributions during inference [23, 24, 25] to improve visual grounding without altering model parameters. However, these methods still incur notable computational overhead due to multi-step procedures such as additional inference under input perturbation or recalibration of attention scores. Consequently, these limitations highlight the need for a streamlined approach that can better utilize visual information during LVLM decoding without additional training, external models, or complex multi-step inference while maintaining computational efficiency.
Figure 1: An overview of ReVisiT. Given an input image and text prompt, the LVLM first encodes the image into vision tokens through a vision encoder and a cross-modal projector. ReVisiT re-purposes these vision tokens as reference informers to guide the text generation process. At each decoding step, ReVisiT (1) constrains the vocabulary $\mathcal{V}$ to $\mathcal{V}_{\texttt{cons}}^{t}$, (2) projects vision token embeddings into $\mathcal{V}_{\texttt{cons}}^{t}$ and selects most relevant token, and (3) refines the final output distribution. Specifically, ReVisiT adoptively leverages vision tokens to serve as reference signals for decoding, enhancing visual grounding without additional forward passes or external supervision.
Contribution. In this work, we propose a simple yet effective decoding method by Referencing Vision tokens to guide the Text generation process of LVLMs (ReVisiT). Vision tokens inherently contain semantic information aligned with textual representations, as they are mapped to the linguistic embedding space through the cross-modality projector. Despite this, they are typically only utilized as the additional input context as same as text tokens, which might be insufficient to use their rich semantic information. Motivated by this, our key insight is to explicitly extract this semantics embedded in vision tokens and leverage them as internal reference signals during decoding. However, directly leveraging these embeddings could be difficult due to their inherent noise and complexity.
To address this challenge, ReVisiT identifies the vision token that best aligns with the current output distribution by minimizing a constrained divergence at each decoding step. Then, ReVisiT uses it as a reference logit to refine the output distribution to better incorporate visual semantics (see Figure 1 for the overview). Specifically, ReVisiT first projects the embeddings of all vision tokens into the text token distribution space. Since vision tokens remain fixed throughout the decoding process, this projection only needs to be computed once before the decoding. The projected probability distributions are constrained to top-$K$ text tokens in the current output distribution, to focus on semantically meaningful ones. ReVisiT then selects the vision token which minimize the Jensen-Shannon Divergence (JSD) between its constrained distribution and the original output distribution, as the most relevant visual guide for each decoding step. Finally, ReVisiT refines the output token distribution by re-weighting it using the constrained distribution of the selected vision token.
We validate the effectiveness of ReVisiT on multiple hallucination benchmarks, including CHAIR [27], POPE [28], and AMBER [29], using LLaVA-1.5-7B [12] and Qwen2.5-VL-7B [13] as the base LVLMs. Across all benchmarks, ReVisiT consistently improves performance over the vanilla greedy decoding baseline, demonstrating enhanced visual grounding capabilities with only a 1-3% increase in inference time. Furthermore, our method shows competitive or superior results relative to state-of-the-art intra-alignment methods [23, 24, 25, 33] with substantially reduced computational costs as these baselines typically incur up to 100% additional computation overhead. These results highlight that appropriate selection of vision tokens enables them to serve as effective informers, enhancing the visual grounding capabilities of LVLMs while offering a practical free lunch.
2 Preliminary and Motivation
We first describe our problem setup formalizing the conventional decoding process of LVLMs in Section 2.1. Then, in Section 2.2, we motivate our approach by identifying the limitations of existing intra-alignment and early exit methods, and propose leveraging vision token representations as internal reference signals for visual grounding. Finally, Section 2.3 provides qualitative analyses demonstrating that vision token hidden states indeed encode meaningful visual semantics, supporting their use during decoding.
2.1 Problem setup
@@
Let’s denote the given input text tokens (e.g., query) and vision tokens as $\mathbf{x}$ and $\mathbf{v}$, respectively. Specifically, the vision tokens $\mathbf{v}$ are generated from the given input image, by forwarding it into vision encoder and cross-modality projector sequentially [12, 13]. Then, our goal is to generate output text tokens $\mathbf{y}$ using LVLM $\mathcal{M}$; following the auto-regressive nature of LLM, output tokens $\mathbf{y}$ (e.g., response) are sequentially generated, i.e., $y_{t}\sim\mathcal{M}(\cdot|\mathbf{v},\mathbf{x},\mathbf{y}_{<t-1})$. Let assume that LVLM $\mathcal{M}$ has $L$ decoding layers and text token vocabulary $\mathcal{V}$, and denote its hidden state of $l$-th layer as $h^{l}$. Then, at time step $t$, the output probability over $\mathcal{V}$ is calculated as below:
$p(h_{T+t-1}^{L},\mathcal{V})=\mathrm{softmax}_{\mathcal{V}}\left(\phi(h_{T+t-1}^{L})\right),~{}\sum_{w\in\mathcal{V}}p(w|h_{T+t-1}^{L},\mathcal{V})=1,$ (1)
@@
$y_{t}\sim p(h_{T+t-1}^{L},\mathcal{V}).$ (2)
2.2 Motivation from prior works
To better align LVLM’s responses with visual semantics in given image, prior works [23, 24, 25] have explored intra-alignment methods. They adjust the model’s output logits (Eq. 2) using the reference logits during decoding. Formally, the decoding process of these methods can be described as:
$y_{t}\sim f\left(p(h_{T+t-1}^{L},\mathcal{V}),\;p(\tilde{h}_{T+t-1}^{L},\mathcal{V})\right),$ (3)
where $\tilde{h}$ denotes the hidden state to obtain the reference logit, and $f$ denotes the logit adjustment operation, such as logit addition, subtraction, or interpolation. For example, VCD [23] adds random noise into input image and forward it to obtain a distorted output logits. M3ID [24] generates vision-agnostic output logits to identify language priors derived from LLM decoder, and SID [25] constructs the reference output logits from low-attention token subsets. While these approaches are effective at aligning visual semantics, they commonly require extra computational cost as reference information is obtained through additional inference passes or recalibration. These observations lead us to the following research question: (RQ) How can we find reference information that is relevant to the visual input to guide decoding and improve visual grounding efficiently?
Interestingly, recent studies in the NLP domain offer a compelling alternative. Rather than relying on externally induced reference signals, the prior works about early exit [34, 35, 36] revealed that intermediate hidden states already encode semantically meaningful information. Specifically, early exit projects intermediate hidden states $h^{j}$ into logits over vocabulary $\mathcal{V}$ using the model’s language modeling head $\phi$, which can be formally expressed as $p(h_{T+t-1}^{j}\mid\mathcal{V})$; here, $j$ denotes the index of an intermediate layer. Motivated by this, DoLa [33] have demonstrated that adjusting logits based on this early exit can effectively resolve output uncertainty during decoding. Roughly, the decoding with early exit can be generalized in the following formula:
$y_{t}\sim f\left(p(h_{T+t-1}^{L},\mathcal{V}),\;p(h_{T+t-1}^{j},\mathcal{V})\right),$ (4)
where the intermediate layer $j$ is selected by maximizing the Jensen-Shannon Divergence (JSD) between the intermediate and final layer logits [33], to identify the layer with the greatest predictive uncertainty. Despite its original design for language-only tasks, DoLa [33] also proves effective for improving visual grounding in LVLMs without any model retraining or architectural changes. These observations indicate that visually relevant auxiliary signals may reside within the model itself, without requiring external augmentation like aforementioned intra-alignment methods.
Prior early exit works have primarily focused on adjusting output distributions based on intermediate representations corresponding to the language modality. Namely, they largely overlook the vision tokens attended by the decoder. However, as LVLMs encode visual information through vision tokens alongside text tokens in the same decoding sequence, intermediate representations of vision tokens may also contain rich visual semantics that are highly relevant for grounding. Moreover, leveraging vision token embeddings offers a distinct advantage: they directly originate from the image input, providing an immediate and explicit pathway to extract visually grounded signals. Therefore, we shift our focus towards vision tokens, investigating whether their intermediate hidden representations can serve as effective internal reference signals for visual grounding by projecting them into logits.
Figure 2: Motivation of ReVisiT. We qualitatively analyzed various vision tokens. Dotted arrows represent vision token projection over specified vocabulary set. For each box, representing text token distribution, we annotated top-5 probable text tokens. Left part illustrate the effectiveness of vocabulary constraint, whereas right part shows the distribution shift during ReVisiT. See Appendix C.1 for a detailed discussion of the underlying values and analysis.
2.3 Deriving vision tokens as reference signals
While early exit [34, 35, 36] has initially explored to accelerate inference, logit-lens [37] employed it to analyze the semantic information encoded in text token hidden states. In this work, we refer logit-lens as the projection of vision tokens’ intermediate representation into the text token distribution space for convenience, although our use case differs from its original intent. Specifically, given a vision token hidden state $h_{i}^{j}$ from layer $j$ and vision token index $i$, we project it into the text token distribution over vocabulary $\mathcal{V}$, which yields $p(h_{i}^{j},\mathcal{V})$.
Under this definition, we first conducted analysis by applying logit-lens to the vision tokens $\mathbf{v}$ at the final decoder layer $L$. As shown in Figure 2 (left), directly applying logit lens to vision tokens over full vocabulary $\mathcal{V}$ results in diffuse and uninformative token distributions. However, manually constraining the vocabulary $\mathcal{V}$ to the subset $\mathcal{V}_{\texttt{env}}$ in Figure 2 effectively reveals the relevant visual semantics, as shown by the top-ranked text tokens in the distribution $p(h_{i}^{j},\mathcal{V}_{\texttt{env}})$. This observation suggests that visual semantics are encoded in visual tokens and can provide reference signals based on that to guide decoding, if one can restrict logit-lens to the proper subset in vocabulary.
While manually defined semantic subsets is considerable as demonstrated in Figure 2, it is not scalable nor adaptive across diverse images and tasks. To resolve this limitation, we draw inspiration by analyzing the inherent decoding dynamics of LVLM with hallucination cases. As illustrated in Figure 2 (right), we observe that hallucinations typically occur when the greedy top-1 prediction diverges from the visual content. However, the top-$5$ candidate tokens at hallucinatory starting point often contain visually grounded alternatives such as "mountain", "landscape" or "painting". This result indicates that the LVLM inherently contains grounded candidates and has a chance to avoid the hallucination, even if the top-1 prediction is incorrect. Thus, at each decoding time step, we construct the constrained vocabulary by selecting the high probable candidates from the vanilla output distribution $p(h_{T+t-1}^{L},\mathcal{V})$. Then, by projecting the vision token representation onto these dynamically selected subsets, we can extract visually grounded reference signals without external inference or heavy computational overhead.
3 ReVisiT: Referencing Vision Tokens to Guide Text Generation
Based on insights in Section 2.3, we propose ReVisiT, a decoding method that adaptively leverages intermediate vision token representations as internal reference signals to guide text generation, enhancing visual grounding with minimal computational overhead. At each decoding step, ReVisiT focuses the decoding process onto a context-relevant vocabulary subset and refines the output by referencing the most relevant vision token. To be specific, ReVisiT consists of the following three main steps: (1) adaptively constraining the vocabulary based on the vanilla output distribution, (2) projecting and selecting vision token hidden states within the subset, and (3) refining the output distribution through element-wise combination and normalization to generate the next token.
3.1 Adaptive context-relevant vocabulary subset construction

### v2

# Revisit What You See: Disclose Language Prior in Vision Tokens for LVLM Decoding

@@ Abstract
Large Vision–Language Models (LVLMs) achieve strong performance across multimodal tasks by integrating visual perception with language understanding. However, how vision information contributes to the model’s decoding process remains under-explored, as reflected in frequent hallucinations. Through a series of analyses, we found that (i) vision tokens provide meaningful visual information even when hallucinations occur, and (ii) their semantics are encoded in the textual space and become explicit under appropriate vocabulary constraints. Building on these observations, we propose ReVisiT, a simple training-free decoding method that references vision tokens to guide text generation. Our approach leverages the semantic information embedded within vision tokens by projecting them into the text token distribution. Specifically, ReVisiT dynamically selects the most relevant vision token at each decoding step via context-aware constrained divergence minimization, and using its constrained projection to refine the output distribution to better incorporate visual semantics. Across five benchmarks on recent LVLMs, ReVisiT consistently enhances visual grounding with minimal computational overhead, and achieves competitive or superior results to state-of-the-art decoding baselines while reducing computational cost by up to $2\times$.  [^1] Code is available at https://github.com/bscho333/ReVisiT.
@@ 1 Introduction
With the recent success of Large Language Models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Team et al., 2023), Large Vision-Language Models (LVLMs) have emerged as powerful multimodal architectures that integrate visual perception with language understanding and text generation (Dai et al., 2023; Liu et al., 2024a; Zhu et al., 2024; Bai et al., 2025; Zhu et al., 2025). Typically, LVLMs encode visual inputs into the LLM decoder’s text embedding space as vision tokens, which are then processed alongside text tokens during decoding. This way of treating vision tokens as static auxiliary contexts, similar to retrieval-augmented generation (Lewis et al., 2020; Gao et al., 2023) in LLMs, enables the construction of complex multimodal systems in a relatively simple manner. However, this approach is often insufficient to capture the unique characteristic and role of vision tokens as sole carriers of visual information, as denoted by hallucinations frequently observed in LVLM’s text outputs (Leng et al., 2024; Favero et al., 2024; Huo et al., 2025; Rohrbach et al., 2018; Li et al., 2023b; Guan et al., 2024). To mitigate this, research has begun to expand the understanding of vision tokens along with their additional utilization (Jiang et al., 2025a; b). Yet, it is still under-explored how vision tokens influence the decoding process of LVLMs and what kinds of textual semantics they encode.
To investigate this direction, we conduct a series of analyses on vision tokens, leading to two key insights. First, by examining the output distributions of LVLMs particularly at hallucinated steps, we find that ground-truth objects often remain among the model’s high-probability candidates and this visual grounding signal in output distribution indeed comes from the vision tokens. Second, we interpret the semantics of vision tokens by projecting them into the text token distribution using the LLM decoder’s language modeling head, as both are embedded in the same space. Here, we observe that the textual semantics of vision tokens is initially obscured, but the interpretable and well-aligned semantics could be discovered in the projected distribution with the proper constraints to focus on specific subset. Overall, these findings suggest that vision tokens intrinsically encode object-level semantics, though such signals typically remain latent under conventional unconstrained decoding.
Building on this observation, we propose ReVisiT, a simple yet effective decoding strategy that Referencing Vision Tokens to guide the text generation process of LVLMs. At each step, ReVisiT selects the most relevant vision token based on context-aware constrained divergence from the current output distribution, and uses its constrained projection as reference to refine token-level logits. This activates the intrinsic semantics of vision tokens without requiring additional training, architectural modifications, external modules, or multi-pass inference.
Figure 1: ReVisiT across various model sizes and architectures. We evaluate on the CHAIR benchmark and report F1. Consistent improvements hold across different size buckets (7–8B / 13–14B / 26–32B) and architectures, demonstrating scalability and model-agnostic effectiveness. Full results are provided in Table 9.
As highlighted in Fig. 1,  ReVisiT consistently improves performance across six model sizes spanning three architecture, demonstrating scalability and model-agnostic effectiveness. To be specific, on three hallucination benchmarks—HallusionBench (Guan et al., 2024), CHAIR (Rohrbach et al., 2018), and POPE (Li et al., 2023b)—ReVisiT reduces hallucinations robustly on both LLaVA-1.5 (Liu et al., 2024a) and Qwen2.5-VL (Bai et al., 2025). Beyond hallucination, we further validate its effectiveness on general-purpose benchmarks, including VQAv2 (Antol et al., 2015) and MMMU (Yue et al., 2024), indicating that the semantics surfaced by our context-aware constrained projection generalize across tasks and architectures.
In summary, our contributions are as follows:
  • $\circ$
    We provide quantitative and qualitative evidences that vision tokens intrinsically encode interpretable semantics exposed when projected over a semantically coherent vocabulary subset.
  • $\circ$
    We introduce ReVisiT, a simple yet effective LVLM decoding method that Referencing Vision Tokens via context-aware constrained divergence.
  • $\circ$
    We validate ReVisiT across six model sizes spanning three architectures, reducing hallucinations and yielding gains on five benchmarks, while maintaining inference cost.
2 Motivation
Figure 2: Motivation. We qualitatively analyzed various vision tokens with LLaVA-1.5-7B (Liu et al., 2024a). Dotted arrows represent vision token projection over specified vocabulary set. For each box, representing text token distribution, we annotated top-5 probable text tokens. Left part illustrate the effectiveness of vocabulary constraint, whereas right part shows the distribution shift during ReVisiT. See Appendix C.1 for a detailed discussion of the underlying values and analysis.
In this section, we investigate whether vision tokens encode interpretable semantics in LVLM’s text token space that can directly inform its decoding process. To conduct systematic analyses, we adopt the CHAIR benchmark (Rohrbach et al., 2018), as it provides a principled definition of hallucinations based on 80 MSCOCO (Lin et al., 2014) object categories with 403 synonyms in total; for each sample, the set of synonyms of annotated object categories is regarded as ground-truth (GT) objects, while the remaining category synonyms are regarded as hallucinated (Hal) objects. We refer to a hallucinated step as a decoding step where generated token corresponds to an object in this MSCOCO synonym set but is not part of the ground-truth annotations for the given image.
Building on this definition, we sample 500 images from the MSCOCO validation set and generate captions for all images. All quantitative analyses are performed with Qwen2.5-VL-7B (Bai et al., 2025) using vanilla greedy decoding with a maximum of 512 tokens, prompted with "Please describe this image in detail.", which is a standard input prompt for CHAIR.
2.1 How do vision tokens affect the output distribution of LVLM?
To examine how vision tokens affect the decoding step of LVLMs, we first ask whether ground-truth (GT) objects remain among the model’s top-probability predictions at hallucinated steps. We use a at-least-one recall@k defined on a probability distribution over the vocabulary: for a given distribution, we measure whether at least one GT object synonym appears within its top-$k$ predictions. Among the 500 generated captions, 176 captions contain at least one hallucinated step, and we obtain 190 hallucinated steps for step-level analysis. For each hallucinated step, we report at-least-one recall@k on the model’s output distribution at that step, excluding the greedy (top-1) prediction from the candidate set (details in Appendix A.1.2).
Figure 3: Ground-truth objects frequently remain in top-probability predictions. At 190 hallucinated steps, at least one GT object is recalled in 95.8% of cases within the top-500 predictions, corresponding to only 0.33% of full vocabulary (151,665 tokens).
The results are summarized in Fig. 3. It is observable that GT objects frequently remain accessible in output distribution: in 63.2% of hallucinated steps, at least one GT object appears within the top-50 candidates, and this recall rises to 95.8% within the top-500. Given the vocabulary size of 151,665 tokens, top-500 corresponds to only 0.33% of the full vocabulary. A qualitative example is also presented in Fig. 2 (right), where the model incorrectly selects “person” as the greedy prediction, yet visually grounded alternatives such as “mountain,” “landscape,” and “painting” appear among the top-5 candidates. This indicates that relevant visual candidates are assigned high probabilities in the output distribution, showing that the LVLM indeed possesses the correct visual information even when hallucinations occur. The 95% confidence intervals are reported in Appendix A.1.2.
We further conduct a fine-grained comparison between the occurrences of GT and hallucinated object subsets during the decoding. Specifically, for each generated token, we track the maximum probability among objects in each subset. Then, we report the average of these per-step maxima as GT Max Avg for GT object subsets and Hal Max Avg for hallucinated object subsets, respectively (see detailed definition in Appendix A.1.3). To investigate the contribution of vision tokens by isolating them, we additionally report w/o image as an image-absent setting where we remove all vision tokens and perform teacher forcing with the previously generated caption as input while keeping decoding otherwise identical.
Table 1: Comparison of Hal Max Avg and GT Max Avg. The results for (i) all output tokens across all 500 CHAIR samples (88,706 tokens), (ii) all output tokens within hallucinated samples (176 samples, 35,116 tokens), and (iii) hallucinated decoding steps within hallucinated samples (176 samples, 190 steps). The higher value between Hal and GT is highlighted in bold.
Analysis Setting Condition GT Max Avg Hal Max Avg
All tokens w/ image 24.49 20.24
w/o image 21.83 22.66
All tokens w/ image 24.42 21.74
(in hallucinated samples) w/o image 21.73 23.88
Hallucinated tokens w/ image 32.00 36.44
w/o image 30.26 35.23
As shown in Table 1, with vision tokens present, GT objects receive higher probability mass than hallucinated ones on average (GT >> Hal); removing the image tokens reverses the ordering (Hal >> GT), confirming a grounding contribution from visual input. At hallucinated steps, the reversed ordering (Hal >> GT) persists even with the presentation of image, indicating that language priors (i.e., textual co-occurrence biases from pre-training (Leng et al., 2024; Favero et al., 2024)) can dominate the prediction at specific positions. We further validate this ordering shift using a paired nonparametric Wilcoxon signed-rank test (Wilcoxon, 1945), which compares the signs and magnitudes of paired step-level differences without assuming normality, showing a significant shift toward GT when moving from w/o-image to w/-image condition in the full-token analyses (Table 1, settings (i) and (ii)), while no significant shift is observed on the hallucinated-step subset (See Appendix A.1.3 for detailed discussion).
2.2 What textual semantics are encoded in vision tokens?
To identify the source of visual grounding signals, we focus on vision tokens, the sole carriers of visual information in LVLMs. Vision tokens share the same embedding space with text tokens and are autoregressively processed during decoding like them. This motivates probing their representations by projecting them through the model’s language-modeling head to obtain the text token distribution over the vocabulary, analogous to logit-lens and early-exit analyses for text token embeddings (nostalgebraist, 2020; Chuang et al., 2024; Wang et al., 2025). Specifically, for each of 500 MSCOCO validation images, we project each vision token’s final-layer hidden state to obtain the text token distribution, yielding 173,173 vision token projections. (see details in Appendix A.1.4).
Figure 4: Vision token projection. Vision tokens reveal rich semantics when projected over semantically coherent subsets.
First, to assess the presence of visual grounding signals, we compute token-level at-least-one recall@k for each vision token projection, following the definition in Sec. 2.1. Quantitatively, across 173,173 vision tokens, at-least-one recall@k is low under the full vocabulary: only 2.03% of projections surface a GT synonym at Top-1, increasing to 21.72% at Top-30. Qualitative inspection also indicates that tokens corresponding to background regions often yield generic function words, with object labels rarely appearing (red boxes on the left side of Fig. 2). These observations indicate that, under the full vocabulary, vision-token projections are diffuse and difficult to interpret.
In contrast, constraining the projection to semantically coherent vocabulary subsets reveals substantially clearer alignment with visual semantics. When restricting the prediction space to the CHAIR-derived object subset (176 single-token synonyms covering 62 categories), at-least-one recall@k on the same 173,173 projections increases sharply to 40.44% at Top-1—about $20\times$ higher than the naïve full-vocabulary projection (2.03%). At Top-30, recall reaches 89.24% while inspecting only $30/176\!\approx\!17\%$ of the subset. For comparability, both settings evaluate GT synonyms from the same 62 categories. Moreover, when we manually define domain-specific subsets, such as environment-related terms including “river,” “mountain,” and “tree”, the projected distribution aligns with visually salient regions and reveals object labels that remain hidden under the full-vocabulary setting (yellow boxes on the left side of Fig. 2).
These findings provide both quantitative and qualitative evidence that vision tokens intrinsically encode object-level semantics, although such signals are not easily observable without appropriate semantic constraints. Details are presented in Appendix A.1.4.
3 ReVisiT: Referencing Vision Tokens to Guide Text Generation
Based on insights from Sec. 2, we propose ReVisiT, a decoding method that adaptively leverages intermediate vision token representations as internal reference signals to guide text generation, enhancing visual grounding with minimal computational overhead. At each decoding step, ReVisiT focuses the decoding process onto a context-relevant vocabulary subset and refines the output by referencing the most relevant vision token. To be specific, ReVisiT consists of the following three main steps: (1) adaptively constraining the vocabulary based on the vanilla output distribution, (2) projecting and selecting vision token hidden states within the subset, and (3) refining the output distribution through element-wise combination and normalization to generate the next token.
@@ 3 ReVisiT: Referencing Vision Tokens to Guide Text Generation
Preliminary. Let denote input text tokens (e.g., query) and vision tokens as $\mathbf{x}$ and $\mathbf{v}$, respectively. Specifically, the vision tokens $\mathbf{v}$ are generated from the given input image, by forwarding it into vision encoder and cross-modality projector sequentially (Liu et al., 2024a; Bai et al., 2025). Then, our goal is to generate output text tokens $\mathbf{y}$ using LVLM $\mathcal{M}$; following the auto-regressive nature of LLM, output tokens $\mathbf{y}$ (e.g., response) are sequentially generated, i.e., $y_{t}\sim\mathcal{M}(\cdot|\mathbf{v},\mathbf{x},\mathbf{y}_{<t-1})$. Let assume that LVLM $\mathcal{M}$ has $L$ decoding layers and text token vocabulary $\mathcal{V}$, and denote its hidden state of $l$-th layer as $h^{l}$. The output distribution is parameterized by model weights $\theta$ and denoted as $p_{\theta}$. Then, at time step $t$, the output probability over $\mathcal{V}$ is calculated as below:
$p_{\theta}(h_{T+t-1}^{L},\mathcal{V})=\mathrm{softmax}_{\mathcal{V}}\left(\phi(h_{T+t-1}^{L})\right),~\sum_{w\in\mathcal{V}}p_{\theta}(w|h_{T+t-1}^{L},\mathcal{V})=1,$ (1)
@@
$y_{t}\sim p_{\theta}(h_{T+t-1}^{L},\mathcal{V}).$ (2)
Figure 5: An overview of ReVisiT. At each decoding step, ReVisiT (1) constrains the vocabulary $\mathcal{V}$ to $\mathcal{V}_{\texttt{cons}}^{t}$, (2) projects vision token embeddings into $\mathcal{V}_{\texttt{cons}}^{t}$ and selects most relevant token, and (3) refines the final output distribution. ReVisiT leverages vision tokens to serve as reference signals for decoding, enhancing visual grounding without additional forward passes or external supervision.
3.1 Adaptive context-aware vocabulary subset construction